1. Consider a newly issued TIPS security with par value $10,000 and 2.6% coupon rate (paid semiannually).

a). If inflation over the next 6 months turns out to be 3.0% (on annualized basis), what is the dollar coupon interest that will be paid in 6 months?

b). If inflation turns out to be 2.0% (annualized) over the following 6 month period, what is the dollar coupon interest that will be paid in 1 year?
Answer #1

a. Inflation adjusted rate = 3% /2 =1.5%

Inflation adjuted principal at the end of 6 months = 10,000 * (1+1.5%) = 10150

Coupon payment = 10150 * 2.6% *1/2 = 131.95

b. Inflation adjusted rate = 2% /2 =1%

Inflation adjuted principal at the end of 6 months = 10150 * (1+1%) = 10251.5

Coupon payment = 10251.5 * 2.6% *1/2 = 133.2695